# Stabilizing black-box model selection with the inflated argmax

### v2

@@ Abstract
Model selection is the process of choosing from a class of candidate models given data. For instance, methods such as the LASSO and sparse identification of nonlinear dynamics (SINDy) formulate model selection as finding a sparse solution to a linear system of equations determined by training data. However, absent strong assumptions, such methods are highly unstable: if a single data point is removed from the training set, a different model may be selected. In this paper, we present a new approach to stabilizing model selection with theoretical stability guarantees that leverages a combination of bagging and an “inflated” argmax operation. Our method selects a small collection of models that all fit the data, and it is stable in that, with high probability, the removal of any training point will result in a collection of selected models that overlaps with the original collection. We illustrate this method in (a) a simulation in which strongly correlated covariates make standard LASSO model selection highly unstable, (b) a Lotka–Volterra model selection problem focused on identifying how competition in an ecosystem influences species’ abundances, and (c) a graph subset selection problem using cell-signaling data from proteomics. In these settings, the proposed method yields stable, compact, and accurate collections of selected models, outperforming a variety of benchmarks.
@@ 1 Introduction
Model selection is typically formulated as a procedure to identify the model that “best” represents data from among a set of candidate models. For instance, scientists may choose among many interpretable models to identify salient variables, interactions, or policies, with the goal of developing a fundamental understanding of the natural or social phenomenon underlying experimental data. Examples of model selection include variable selection (choosing a subset of covariates from a larger pool that best explains the response variable or label), selecting a decision tree, selecting the order or memory of an autoregressive model, selecting a kernel function for a support vector machine, selecting the number of clusters in k-means, selecting the number of factors in factor analysis, and more. The issue of model selection is a common problem in a variety of domains, including bioinformatics (Saeys et al.,, 2007), environmental studies (Effrosynidis and Arampatzis,, 2021), and psychology (Vrieze,, 2012).
@@ 1.1 Our contributions
We show empirically that the inflated argmax can improve model selection stability, and for variable selection (a special case of model selection), we similarly show this result for data with correlated covariates, a setting known to yield instabilities in standard model selection algorithms. We compare against and outperform numerous conventional model selection approaches, described in § 2.2. These conventional methods, unlike our approach, are not adaptive to the underlying level of uncertainty of the model selection procedure.
@@ 2 Model selection
More formally, model selection focuses on selecting one or more models from among a set of candidate models, denoted $M^{+}$. A model selection procedure is a map $\mathcal{M}$ from data $\mathcal{D}$ to a set of selected models $\hat{M}\subseteq M^{+}$. The output $\hat{M}=\mathcal{M}(\mathcal{D})$ is a nonempty subset of $M^{+}$. If $\mathcal{M}$ always a singleton (i.e., $|\hat{M}|=1$), $\mathcal{M}$ is called simple.
Example. In variable selection (a special case of model selection), the data consists of $n$ variable-response pairs $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$. We seek to identify relevant variables, so the model class $M^{+}=2^{[d]}$ is the power set of $[d]=\{1,\ldots,d\}$.
The set of all candidate models $M^{+}$ can contain any class of models. For instance, in graphical model selection, $M^{+}$ is the set of all graphs $G$ on $d$ vertices, where an edge $(j,k)$ in $G$ encodes conditional dependence between variables $j$ and $k$ (Drton and Perlman,, 2004; Friedman et al.,, 2008). In dynamical system identification, $M^{+}$ is a set of possible differential equations governing the dynamics underlying noisy data (Brunton et al.,, 2016). In econometrics, $M^{+}$ could be a collection of competing time series forecasts (Masini et al.,, 2023). To keep our discussion general, the model class is an abstract, countable set $M^{+}$.
@@
where $\mathcal{A}$ maps data $\mathcal{D}$ to weights $\hat{w}=\mathcal{A}(\mathcal{D})\in\mathbb{R}^{|M^{+}|}$, where for each $m\in M^{+}$, $\hat{w}_{m}$ is the weight assigned to model $m\in M^{+}$, and this weight can be thought of as a probability of each candidate model. The selection rule $\mathcal{S}$ takes these weights $\hat{w}$ as input and returns a non-empty set of selected model(s) $\hat{M}\subseteq M^{+}$ that provide the best fit to the data.
@@ 2.1 Stage 1: Assigning weights to candidate models with $\mathcal{A}$
Bagged algorithm $\tilde{\mathcal{A}}_{K,B}$. We denote bagging a base algorithm $\mathcal{A}$ as $\tilde{\mathcal{A}}_{K,B}$, where $K$ is the number of samples per bag, and $B$ is the number of bags. The output of the weighting algorithm $\tilde{\mathcal{A}}_{K,B}$ must be bounded. For simplicity, we assume that the weights can be normalized such that $\hat{w}$ belongs to the probability simplex $\Delta_{|M^{+}|-1}$.  [^1] The probability simplex $\Delta_{|M^{+}|-1}$ consists of nonnegative weights $w$ that sum to one: $\sum_{m\in M^{+}}w_{m}=1$. Bagging averages an ensemble of weight vectors fit on different randomly sampled bags — see Algorithm 1. Sampling with replacement is known as bagging, and sampling without replacement is known as subbagging.
For simple model selection procedures, the ensemble weight $\hat{w}_{m}$ simply counts the fraction of bags wherein model $m$ was selected. Even if the set of candidate models is infinite (i.e., $|M^{+}|=\infty$), the weights $\hat{w}$ will have at most $B$ nonzero entries, so bagging is still tractable. Generally, increasing the number of bags $B$ used in $\tilde{\mathcal{A}}_{K,B}$ provides a more stability, as we empirically show in Figures 12 and 13.
@@ 2.2 Stage 2: Conventional selection rules $\mathcal{S}$
We discuss conventional selection approches $\mathcal{S}$ and their limitations.
@@
Stability selection. In the context of variable selection, the stability selection method (Meinshausen and Bühlmann,, 2010) selects a single model based on whether each covariate is included in the model with probability at least $\tau$. Concretely, consider a vector of weights $\hat{w}$ indexed by $M^{+}=2^{[d]}$, where $\hat{w}_{m}$ roughly represents the probability that $m$ is the correct model. Define the marginal variable probabilities $\hat{\Pi}_{j}=\sum_{\begin{subarray}{c}m\subseteq[d]:j\in m\end{subarray}}\hat{w}_{m},$ for $j=1,\ldots,d,$ where $j\in m$ corresponds to variable $j$ being included in model $m$. Stability selection returns
$\text{ip}_{\tau}(\hat{w}):=\Big{\{}j\in[d]:\hat{\Pi}_{j}\geq\tau\Big{\}},$
for some use-specified threshold $\tau$. Similar to the standard argmax (without ties), $\text{ip}_{\tau}(\hat{w})$ only returns a collection of variables, which together constitute a single model. This selection procedure can be unstable if some marginal probabilities $\hat{\Pi}_{j}$ are near the threshold $\tau$. It also may fail to adapt to relationships among variables (see Appendix A for an example).
@@ 2.3 Model selection stability
To assess the stability of $\mathcal{M}$, we measure how the output of $\mathcal{M}$ changes with small perturbations to $\mathcal{D}$. We adapt the definition of stability in Soloff et al., 2024b to the context of model selection.
@@
In Definition 2.1, $\mathcal{D}^{\setminus i}$ refers to dataset $\mathcal{D}$ with the $i$-th sample removed. In plain language, the inequality in (2) quantifies stability in terms of whether the sets of selected models overlap when we drop a single observation at random. Note that this notion of model selection stability holds for any “black box” model selection method of the form (1), in contrast to prior works that focus exclusively on variable selection (e.g., Meinshausen and Bühlmann, (2010)).
@@ 3 Related work
Returning a model selection set. Returning a set of models to acknowledge inherent uncertainty in model selection is an old idea. Early literature on subset selection for linear regression (Spjøtvoll,, 1972) focused on selecting the true model with high probability. More recent works extend the idea of a model confidence set to more general modeling settings (Hansen et al.,, 2011; Zhang et al.,, 2024). These works have a stochastic model for the data and construct a confidence set containing the best-fitting model with high probability. By contrast, our framework places no stochastic assumptions on the data, instead targeting a more modest goal of stability for the selected set.
Stability in variable selection. There is extensive literature on stability in variable selection; see, e.g., Khaire and Dhanalakshmi, (2022) for a recent overview. Work in this area has overwhelmingly focused on “simple” procedures, returning a single subset of variables, and stability is measured by the degree of similarity of the selected subsets of variables. By contrast, Definition 2.1 is not satisfied by simple procedures unless they return the same model most of the time.
@@
Stabilizing classification. Our work builds on ideas from stable classification. Soloff et al., 2024b introduces a framework for set-valued classification, combining the inflated argmax with bagging to guarantee stability for any base model selection method.
@@ 4.1 The inflated argmax for model selection
In this section, we propose an alternative selection procedure $\mathcal{S}$ based on the inflated argmax. Soloff et al., 2024b introduced the inflated argmax to guarantee stability in the context of multiclass classification. A key insight of this paper is that if we formulate model selection as a multi-class classification problem, where each class corresponds to one of the candidate models, then we can directly use the inflated argmax and inherit its corresponding theoretical stability guarantees for model selection. With this insight, the inflated argmax, an alternative model selection criteria $\mathcal{S}$ that we denotd by argmaxε, will allow us to decide which subset $\hat{M}$ of $M^{+}$ to return.
@@
where dist$(w,R_{m}^{\varepsilon})=\inf_{q\in R_{m}^{\varepsilon}}||q-w||.$
A visualization of a simple example of argmaxϵ where more than one model is returned is shown in Figure 1. An intuitive explanation of Equation 4 is that $m\in\text{argmax}^{\varepsilon}(w)$ indicates that $m$ would have the largest weight in $w$ by some margin if the weights $w$ were slightly perturbed, where both the size of the margin and the weight perturbations scale with $\varepsilon$. This definition naturally allows for the number of returned models to adapt to the uncertainty in the model selection: for example, if the top two competing models are close in probability, then the inflated argmax will return both models for some choice of small $\varepsilon$. On the other hand, if the top two competing models have a large enough gap in their two weights, the inflated argmax will only return the top model for a choice of small $\varepsilon$.
The parameter $\varepsilon$ is not something to be tuned. It is chosen to satisfy a user-specified stability tolerance; see § 4.2.
@@ 4.2 Stability guarantee
In this section, state a theorem guaranteeing the stability of our approach based on our definition of stability in Definition 2.1.
@@
The main use for this theorem is that the user can choose a tolerable worst-case instability $\delta$ for their model selection, input the known parameters of the problem (i.e., $|M^{+}|$ candidate models, $K$ samples per bag, $n$ total samples, and $B$ bags), and directly solve for the $\varepsilon$ that should be used for argmaxε.
Recall that $\varepsilon$ determines how much we inflate the argmax, so when it’s smaller, the inflated argmax is closer to the standard argmax. Accordingly, when $\varepsilon$ is smaller, Theorem 1 gives a larger $\delta$, reflecting a weaker stability guarantee. Further, note that the instability level $\delta$ decreases (i.e., stability increases) by decreasing the bag size $K$ and increasing the number $B$ of bags. The size of our model class $|M^{+}|$ has only a small influence on the level of instability of our procedure. For example, when there are two models, $|M^{+}|=2$, the factor $1-\frac{1}{|M^{+}|}=\frac{1}{2}$, giving a constant-factor boost in stability over the case where $M^{+}$ is countably infinite, where $1-\frac{1}{|M^{+}|}$ becomes $1$.
This theoretical result is the first to provide theoretical guarantees for model selection stability that is agnostic to the underlying models, base algorithm, and data distribution.
5 Experiments
Our experiments assess and compare the stability of five different model selection approaches, described in Table 1.
@@ 5 Experiments
$\mathop{\!\arg\!\max}$ $\circ\ \mathcal{A}$ Model(s) with maximum weight assigned by $\mathcal{A}$
@@
We perform experiments in three settings: identifying important variables in a synthetic dataset that contains a small collection of highly correlated variables (see Appendix C), identifying the latent Lotka-Volterra governing differential equations from noisy trajectory data generated from these equations, and constructing a graph to represent cell-signaling pathways of 11 proteins. In the first two settings, we generate synthetic datasets for $N$ trials, each contaminated with independent Gaussian noise, allowing us to examine the stability across a collection of datasets where we know the ground truth $m^{*}$.

### v3

@@ Abstract
Model selection is the process of choosing from a class of candidate models given data. For instance, methods such as the LASSO and sparse identification of nonlinear dynamics (SINDy) formulate model selection as finding a sparse solution to a linear system of equations determined by training data. However, absent strong assumptions, such methods are highly unstable: if a single data point is removed from the training set, a different model may be selected. In this paper, we present a new approach to stabilizing model selection with theoretical stability guarantees that leverages a combination of bagging and an “inflated” argmax operation. Our method selects a small collection of models that all fit the data, and it is stable in that, with high probability, the removal of any training point will result in a collection of selected models that overlaps with the original collection. We illustrate this method in (a) a simulation in which strongly correlated covariates make standard LASSO model selection highly unstable, (b) a Lotka–Volterra model selection problem focused on identifying how competition in an ecosystem influences species’ abundances, (c) a graph subset selection problem using cell-signaling data from proteomics, and (d) unsupervised $\kappa$-means clustering. In these settings, the proposed method yields stable, compact, and accurate collections of selected models, outperforming a variety of benchmarks.
@@ 1 Introduction
Model selection is typically formulated as a procedure to identify the model that “best” represents data from among a set of candidate models. For instance, scientists may choose among many interpretable models to identify salient variables, interactions, or policies, with the goal of developing a fundamental understanding of the natural or social phenomenon underlying experimental data. Examples of model selection include variable selection (choosing a subset of covariates from a larger pool that best explains the response variable or label), selecting a decision tree, selecting the order or memory of an autoregressive model, selecting a kernel function for a support vector machine, selecting the number of clusters in $\kappa$-means, selecting the number of factors in factor analysis, and more. The issue of model selection is a common problem in a variety of domains, including bioinformatics (Saeys et al.,, 2007), environmental studies (Effrosynidis and Arampatzis,, 2021), and psychology (Vrieze,, 2012).
@@ 1.1 Our contributions
We show empirically that the inflated argmax can improve black-box model selection stability with examples in clustering, decision trees, and variants of variable selection tasks. For variable selection (a special case of model selection), we show this result for data with correlated covariates, a setting known to yield instabilities in standard model selection algorithms. We compare against and outperform numerous conventional model selection approaches, described in § 2.2. These conventional methods, unlike our approach, are not adaptive to the underlying level of uncertainty of the model selection procedure.
@@ 2 Model selection
More formally, model selection focuses on selecting one or more models from among a set of candidate models, denoted $M^{+}$. A model selection procedure is a map $\mathcal{M}$ from data $\mathcal{D}$ to a set of selected models $\hat{M}\subseteq M^{+}$. [^1] Note that this is a general formulation of the model selection problem, and does not assume that there is a true data generating model in $M^{+}$. The output $\hat{M}=\mathcal{M}(\mathcal{D})$ is a nonempty subset of $M^{+}$. If $\mathcal{M}$ always outputs a singleton (i.e., $|\hat{M}|=1$), $\mathcal{M}$ is called simple.
For example, in variable selection (a special case of model selection), the data consists of $n$ variable-response pairs $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$. We seek to identify relevant variables (e.g., variables of $x$ most predictive of $y$.), so the model class $M^{+}=2^{[d]}$ is the power set of $d]=\{1,\ldots,d\}$. However, in more general settings, the set of all candidate models $M^{+}$ can contain any class of models. For instance, in graphical model selection, $M^{+}$ is the set of all graphs $G$ on $d$ vertices, where an edge $(j,k)$ in $G$ encodes conditional dependence between variables $j$ and $k$ (Drton and Perlman,, 2004; Friedman et al.,, 2008). In dynamical system identification, $M^{+}$ is a set of possible differential equations governing the dynamics underlying noisy data (Brunton et al.,, 2016). In econometrics, $M^{+}$ could be a collection of competing time series forecasts (Masini et al.,, 2023). To keep our discussion general, the model class is an abstract, countable set $M^{+}$.
@@
where $\mathcal{A}$ maps data $\mathcal{D}$ to weights $\hat{w}=\mathcal{A}(\mathcal{D})\in\mathbb{R}^{|M^{+}|}$, where for each $m\in M^{+}$, $\hat{w}_{m}$ is the weight assigned to model $m\in M^{+}$. This weight may be loosely interpreted as the probability of the candidate model being the best model given the data among the choices in $M^{+}$. The selection rule $\mathcal{S}$ takes these weights $\hat{w}$ as input and returns a non-empty set of selected model(s) $\hat{M}\subseteq M^{+}$ that provide the best performance as a function of the training data, where performance may reflect some combination of fit to data and structural properties of the model.
@@ 2.1 Stage 1: Assigning weights to candidate models with $\mathcal{A}$
Bagged algorithm $\tilde{\mathcal{A}}_{K,B}$. We denote $\tilde{\mathcal{A}}_{K,B}$ as the algorithm resulting from applying a bagging technique on a base algorithm $\mathcal{A}$, where $K$ is the number of samples per bag and $B$ is the number of bags. The output of the weighting algorithm $\tilde{\mathcal{A}}_{K,B}$ must be bounded. For simplicity, we assume that the weights can be normalized such that $\hat{w}$ belongs to the probability simplex $\Delta_{|M^{+}|-1}$.  [^2] The probability simplex $\Delta_{|M^{+}|-1}$ consists of nonnegative weights $w$ that sum to one: $\sum_{m\in M^{+}}w_{m}=1$. Bagging averages an ensemble of weight vectors fit on different randomly sampled bags; see Algorithm 1. Sampling with replacement is known as bagging, and sampling without replacement is known as subbagging.
For a base algorithm $\mathcal{A}$ that places all weight on a single model, bagging $\mathcal{A}$ to compute $\hat{w}_{m}$ simply counts the fraction of bags where $m$ was selected. Even if the set of candidate models is infinite (i.e., $|M^{+}|=\infty$), the weights $\hat{w}$ will have at most $B$ nonzero entries, so bagging is still tractable. Generally, increasing the number of bags $B$ used in $\tilde{\mathcal{A}}_{K,B}$ provides a more stability, as we empirically show in Figures 12 and 13.
@@ 2.2 Stage 2: Conventional selection rules $\mathcal{S}$
We discuss conventional selection approaches $\mathcal{S}$ and their limitations.
@@
[…]
@@ 2.3 Model selection stability
The stability of $\mathcal{M}$ corresponds to how much the output of $\mathcal{M}$ may change with small perturbations to $\mathcal{D}$ for any dataset $\mathcal{D}$. Note that this cannot be estimated empirically because we cannot test all possible datasets $\mathcal{D}$ with finite computational resources. We adapt the definition of stability in Soloff et al., 2024b to the context of model selection.
@@
In Definition 2.1, $\mathcal{D}^{\setminus i}$ refers to dataset $\mathcal{D}$ with the $i$-th sample removed. In plain language, the inequality in (2) quantifies stability in terms of whether the sets of selected models overlap when we drop a single observation at random. Note that this notion of model selection stability holds for any “black box” model selection method of the form (1), in contrast to prior works that focus on particular settings, such as variable selection (e.g., Meinshausen and Bühlmann, (2010)) or $\kappa$-means clustering (e.g., Ben-David et al., (2007)). While this definition of stability may seem weak (i.e., in practice, much larger perturbations to the data could be made), we will see empirically in our results that this minimal change to the data can have a substantial impact on model selection stability. Moreover, stability according to this criterion is necessary for most stronger notions of stability to be satisfied.
@@ 3 Related work
Model selection has a rich history in the statistics literature. Ding et al., (2018) provides a recent overview of model selection techniques. Various aspects of model selection have been studied in previous work, including hyperparameter selection via cross-validation (Raschka,, 2020), preventing overfitting (Cawley and Talbot,, 2007), selection criteria (Rao et al.,, 2001), among others. In this work, we are particularly interested in the aspect of model selection stability, and bagging has been a key tool to address instability in model selection.
Returning a model selection set. Returning a set of models to acknowledge inherent uncertainty in model selection is an old idea. Early literature on subset selection for linear regression (Spjøtvoll,, 1972) focused on selecting the true model with high probability. Breiman, (2001) coined the terms “Rashomon set” to refer to a complete set of nearly equally performing models (e.g., Xin et al., (2022) finds the full Rashomon set for decision trees), and the “Rashomon effect” to refer to the phenomenon that multiple, vastly distinct models may all perform equally well on out-of-sample data for a particular problem. More recent works extend the idea of a model confidence set to more general modeling settings (Hansen et al.,, 2011; Zhang et al.,, 2024). These works have a stochastic model for the data and construct a confidence set containing the best-fitting model with high probability. By contrast, our framework places no stochastic assumptions on the data, instead targeting a more modest goal of stability for the selected set.
Stability in variable selection. There is extensive literature on stability in variable selection; see, e.g., Khaire and Dhanalakshmi, (2022) for a recent overview. Work in this area has overwhelmingly focused on “simple” procedures, returning a single subset of variables, and stability is measured based on the similarity of the selected subsets of variables. By contrast, Definition 2.1 is not satisfied by simple procedures unless they return the same model most of the time.
@@
Stabilizing classification. Our work builds on ideas from stable classification. Soloff et al., 2024b introduces a framework for set-valued classification, combining the inflated argmax with bagging to guarantee stability for classification.
@@ 4.1 The inflated argmax for model selection
In this section, we propose an alternative selection procedure $\mathcal{S}$ based on the inflated argmax. Soloff et al., 2024b introduced the inflated argmax to guarantee stability in the context of multiclass classification. A key insight of this paper is that if we formulate model selection as a multi-class classification problem, where each class corresponds to one of the candidate models, then we can directly use the inflated argmax and inherit its corresponding theoretical stability guarantees for model selection. With this insight, the inflated argmax, an alternative model selection criteria $\mathcal{S}$ that we denote by argmaxε, will allow us to decide which subset $\hat{M}$ of $M^{+}$ to return.
@@
[…]
A visualization of a simple example of argmaxϵ where more than one model is returned is shown in Figure 1. An intuitive explanation of Equation 4 is that $m\in\text{argmax}^{\varepsilon}(w)$ indicates that $m$ would have the largest weight in $w$ by some margin if the weights $w$ were slightly perturbed, where both the size of the margin and the weight perturbations scale with $\varepsilon$. This definition naturally allows for the number of returned models to adapt to the uncertainty in the model selection: for example, if the top two competing models are close in probability, then the inflated argmax will return both models for some choice of small $\varepsilon$. On the other hand, if the top two competing models have a large enough gap in their two weights, the inflated argmax will only return the top model for a choice of small $\varepsilon$. The parameter $\varepsilon$ is not something to be tuned. It is chosen to satisfy a user-specified stability tolerance; see § 4.2.
@@ 4.2 Stability guarantee
In this section, we state a theorem guaranteeing the stability of our approach based on our definition of stability in Definition 2.1.
@@
The proof immediately follows from the proof presented in Soloff et al., 2024b . The main use for this theorem is that the user can choose a tolerable worst-case instability $\delta$ for their model selection, input the known parameters of the problem (i.e., $|M^{+}|$ candidate models, $K$ samples per bag, $n$ total samples, and $B$ bags), and directly solve for the $\varepsilon$ that should be used for argmaxε.
Recall that $\varepsilon$ determines how much we inflate the argmax, so when it’s smaller, the inflated argmax is closer to the standard argmax. Accordingly, when $\varepsilon$ is smaller, Theorem 1 gives a larger $\delta$, reflecting a weaker stability guarantee. Further, note that the instability level $\delta$ decreases (i.e., stability increases) by decreasing the bag size $K$ and increasing the number $B$ of bags. The size of our model class $|M^{+}|$ has only a small influence on the level of instability of our procedure. For example, when there are two models, $|M^{+}|=2$, the factor $1-\frac{1}{|M^{+}|}=\frac{1}{2}$, giving a constant-factor boost in stability over the case where $M^{+}$ is countably infinite, where $1-\frac{1}{|M^{+}|}$ becomes $1$. This theoretical result is the first to provide theoretical guarantees for model selection stability that is agnostic to the underlying models, base algorithm, and data distribution.
@@
$\mathop{\!\arg\!\max}$ $\circ\ \mathcal{A}\ \$Model(s) with maximum weight assigned by $\mathcal{A}$
@@
5 Experiments
Our experiments assess and compare the stability of five different model selection approaches, described in Table 1. We perform experiments in three settings: identifying important variables in a synthetic dataset that contains a small collection of highly correlated variables (see Appendix C), identifying the latent Lotka-Volterra governing differential equations from noisy trajectory data generated from these equations, and constructing a graph to represent cell-signaling pathways of 11 proteins. In the first two settings, we generate synthetic datasets for $N$ trials, each contaminated with independent Gaussian noise, allowing us to examine the stability across a collection of datasets where we know the ground truth $m^{*}$. The code used to produce the results in this paper can be found at https://github.com/Melissa3248/stable-model-selection.